In Newton's ring experiment the diameter of bright ring is reduced to 0.8 of its value after introducing a liquid below the convex surface.calculate the refractive index at liquid.
1
Expert's answer
2019-08-19T09:41:33-0400

We can write for radius of bright ring

"r_{k}=\frac{\sqrt{(2k-1)Rλ}}{\sqrt{2n}} (1)"


In our two cases

"r_{1}=\frac{\sqrt{(2k-1)Rλ}}{\sqrt{2}} (2)"

"r_{2}=\frac{\sqrt{(2k-1)Rλ}}{\sqrt{2n}} (3)"


According to the condition of the problem, we have

"r_{2}=0.8 r_{1} (4)"



We put (2) and (3) in (4)

n=1.5625


Answer

1.5625